Science Report – May 1st

Crew 193 – PHEnOM Gold Crew

Submitted by XO/Science Officer Doug Campbell

1. Emergency EVAC EVA: Two EVAs took place today; however, no new sites were noted as useful for emergency evacuation shelters.

2. Sunspot and Solar Flare Monitoring: One small, shallow prominence was photographed on the Sun. No other features were apparent.

3. Shortwave Texting on EVA’s: Beartooth devices are still not working. Troubleshooting through contact with the manufacturer is occurring.

4. GPS Route Measurements: Both EVAs took the device with them to map their route. Measurements will be used by a PHEnOM colleague to map the routes taken and determine if the easiest routes were used.

5. Waterless dish cleaning: Bacteria growth and monitoring will continue over the remaining days of the mission. The study is starting to show promising results regarding the use of UV and a cloth to clean dishes on mars without the use of water. Using dirt to clean the dishes seems to have some rapid growing bacteria!

6. The MAG (Maximum Absorbency Garment) utilization study: The study has been scrapped. Results will be from blue crew only.

7. Spacesuit Helmet Fogging: Study has been discontinued after 7 days of data collection. No fogging was ever reported. One helmet fogged today on EVA during a brief rainfall. Once rain stopped, it returned to un-fogged almost immediately.

8. Crew Comfort, Health and Safety on EVA: Measurements of the crew’s weight, blood pressure, SpO2 and pulse were taken before and after the EVA. Maximum heart rate was tracked by a mobile monitor during the EVA when the astronaut felt that they were working hard (normally climbing a large hill. Analysis will be done as the mission continues.

Astronomy Report – May 1st

Astronomy Report

Name: Eric Shear

Crew: 193
Date: May 1, 2018

Sky Conditions: Mostly clear, occasional clouds

Wind Conditions: Low

Observation Start Time: 1330

Observation End Time: 1400


One small, shallow prominence was photographed on the Sun. No other features were apparent.

Photos were taken with the Skyris camera and the best images processed in Autostakkert, Registax, Photoshop, and PowerPoint to colorize and bring out contrast. The results are attached.

Objects Viewed: Sun

Problems Encountered:

One filament was spotted, but was nowhere to be found when viewed through the camera. No amount of tuning could bring it back. The prominence discovered today seems to have developed as quickly.

Greenhab Report – May 1st

GreenHab Report

Date: 1May2018

Environmental control: Ambient with door open

Shade cloth: on

Fan: not used today

Average temperatures: 22.1 ℃

Low: 16.8℃

High: 52.7℃

Hours of supplemental light: None

Changes to crops: Yellowing resolving, plants seem to be in slightly better health.

Daily water usage for crops: 7.75 Gallons

Time(s) of watering for crops: Twice: In the morning at 0800 and at 1730.

Morning research observations: None

Changes to research plants: None

Aquaponics: None

Narrative: Yellowed plants are are turning green again after extra watering. Started singing to them today. Selections included Queen (Greatest Hits). If they rise from the dead overnight, we’ll be on to something.

Support/supplies needed: None

Copyright © The Mars Society. All rights reserved. | Main Site